OpenStudy (anonymous):

simplify\[(\sqrt{-3}+4)(\sqrt{-3}-1)\]

OpenStudy (anonymous):

im coming up with 3-1square root -3 +4square root-3-4 answer -3+4square root 3

OpenStudy (anonymous):

so would it be just square root of 3

OpenStudy (anonymous):

2i^2

OpenStudy (turingtest):

nevermind, misread the Q

OpenStudy (anonymous):

phemoo1 how did you come up with that i must have missed leaving a neg in root to pull out the i

OpenStudy (turingtest):

\[(\sqrt{-3}-4)(\sqrt{-3}+1)\]right?

OpenStudy (anonymous):

yes how did you write that on here would be a huge start and help alot

OpenStudy (turingtest):

use the "equation" feature at the bottom of the little window you have it won't work that well for the original post though, you need to learn the language LaTeX for that I'll do it for you, but experiment with the equation button

OpenStudy (anonymous):

i can do it on the button below but that doesnt post it the first time

OpenStudy (turingtest):

yeah I know, that's why I did it for you anyway, gotta FOIL it out\[(\sqrt{-3}-4)(\sqrt{-3}+1)=\sqrt{-3}\cdot\sqrt{-3}+\sqrt{-3}-4\sqrt{-3}-4(1)\]

OpenStudy (anonymous):

which i did

OpenStudy (anonymous):

whats \[\sqrt{?}\]

OpenStudy (turingtest):

you mean how do you make the symbol?

OpenStudy (anonymous):

3-1\[3-1\sqrt{?}\]

OpenStudy (turingtest):

is that what you get for the simplification? I am confused what you are writing

OpenStudy (anonymous):

\[3-1 \sqrt}-3 +4\sqrt{-3-4\]

OpenStudy (turingtest):

it would be like this if you want to do it the right way...

OpenStudy (turingtest):

\[(\sqrt{-3}-4)(\sqrt{-3}+1)\.] write this without the little dot at the end

OpenStudy (turingtest):

I keep dropping that 4 :/

OpenStudy (turingtest):

take out the dot and you get\[(\sqrt{-3}-4)(\sqrt{-3}+1)\]

OpenStudy (anonymous):

when i foiled it i got \[3-1\sqrt{?}\]

OpenStudy (anonymous):

forget it

OpenStudy (anonymous):

what did you get after foil will tell you if i got same

OpenStudy (turingtest):

after FOILing you get\[\sqrt{-3}\cdot\sqrt{-3}+(1)\sqrt{-3}-4\sqrt{-3}-4(1)\]

OpenStudy (anonymous):

why -4

OpenStudy (anonymous):

and 1 shouldnt it be -1 and +4

OpenStudy (turingtest):